### 12 Dec 2023 v1

@@ Abstract
Uncertainty quantification is a pivotal field that contributes to the realization of reliable and robust systems. By providing complementary information, it becomes instrumental in fortifying safe decisions, particularly within high-risk applications. Nevertheless, a comprehensive understanding of the advantages and limitations inherent in various methods within the medical imaging field necessitates further research coupled with in-depth analysis. In this paper, we explore Conformal Prediction, an emerging distribution-free uncertainty quantification technique, along with Monte Carlo Dropout and Evidential Deep Learning methods. Our comprehensive experiments provide a comparative performance analysis for skin lesion classification tasks across the three quantification methods. Furthermore, We present insights into the effectiveness of each method in handling Out-of-Distribution samples from domain-shifted datasets. Based on our experimental findings, our conclusion highlights the robustness and consistent performance of conformal prediction across diverse conditions. This positions it as the preferred choice for decision-making in safety-critical applications.
@@ 1 Introduction
In machine learning, two sources of uncertainties collectively influence model predictions [7]. The first source, known as epistemic uncertainty, arises from a lack of knowledge and understanding. Model retraining and incorporating additional information and data are typically employed to mitigate this type of uncertainty [8]. The second source of uncertainty is aleatoric uncertainty, which characterizes the inherent randomness within the data and, as such, remains irreducible. In the existing literature, various approaches offer both theoretical and practical solutions for quantifying uncertainties associated with deep learning models. These methods can be broadly categorized into Single Deterministic Networks, Bayesian Neural Networks, Deep Ensembles, and Test-time Augmentation methods [9].
@@
Single deterministic networks are straightforward methods to represent the predictive uncertainty of deep learning models. As suggested by the name, the networks are trained to adjust their weights in a deterministic manner, and the uncertainty is measured through a single forward pass. A simple baseline of single deterministic networks is the utilization of the neural network output probabilities (i.e., SoftMax outputs for classification) as an interpretation of the model uncertainty. Several studies, however, showed that neural networks are often over-confident [10], hence their output probabilities are miscalibrated [11] and tend to generate wrong uncertainty estimates. A more generic approach in single deterministic networks includes predicting the parameters of a prior distribution over the outputs. such algorithms are known as Evidential learning [12] or Prior Networks [13]. The framework of these approaches involves introducing a higher-order distribution over the likelihood function, with the parameters of these distributions being estimated using the underlying neural network. The predicted class probability corresponds to the expected value of the distribution, while the variance of the distribution serves as an estimate of the model’s uncertainty.
@@
Bayesian Neural Networks (BNNs) [14] seamlessly merge the principles of Bayesian inference with traditional neural networks. Instead of considering BNN parameters as fixed, deterministic values, they are treated as probability distributions. This unique approach empowers BNNs to effectively capture and represent predictive uncertainty. Consider a neural network with parameters $\omega$ and a training dataset $D_{t}=(x,y)$. If the network weights are assigned a prior distribution, denoted as $P(\omega)$, the posterior distribution $P(\omega|x,y)$ can be calculated using Bayes’ theorem. To calculate the posterior $p(\omega|X,Y)$, it necessitates integrating over all possible values of $\omega$, a task that is often intractable. Generally, obtaining the posterior is only viable through either Variational Inference [15], where a simple and tractable distribution is used to approximate the exact posterior, or through other approximation techniques, such as Markov Chain Monte Carlo (MCMC) [16].
@@
Recently, Conformal prediction (CP) [18, 19] emerged as a prominent uncertainty quantification technique. It has gained attention in the fields of machine learning and computer vision [20, 21]. In addition to its elegance and computational efficiency, CP algorithms belong to the domain of ’distribution-free’ methodologies. This implies that they are model-agnostic, data distribution agnostic, and applicable to finite sample sizes.
In this paper, we compare three top-performing uncertainty quantification methods, namely CP, Monte Carlo Dropout (MCD), and Evidential Deep Learning (EDL), and provide a comprehensive and detailed analysis of their performance. We include three experiments to underscore the unique attributes of each method, with a particular emphasis on CP, a less-explored approach for skin lesion classification tasks. The main contributions of our work are as follows:
@@
2 Background on Uncertainty quantification Methods
In the rich domain of uncertainty quantification, understanding the nuances of existing methodologies and their implementation details is paramount. This section delves into the background literature and related work that forms the foundation for uncertainty methods, We specifically focus on three quantification techniques that form the basis of our experimental analysis. For a more comprehensive review of uncertainty quantification methods in deep learning, interested readers can refer to [9, 22, 23].
2.1 Monte Carlo Dropout
@@
In the case of N multiple passes, Let $f(x)$ be the prediction of the neural network for input $x$. During MCD, we obtain $T$ samples of predictions $f_{t}(x)$ where $t=1,2,...,T$ represents each Monte Carlo sample. The mean prediction can be represented as:
@@
2.2 Evidential Deep Learning
In a traditional Convolution Neural Network (CNN), label probabilities are generated by applying a SoftMax activation function to the last layer. The objective is to maximize the likelihood function by minimizing a metric of loss, often a cross-entropy loss, which tends to make the prediction "overconfident" [25]. The output of the SoftMax operation reflects discrete point estimates of class probabilities without proper indication of the uncertainty present in the estimate. EDL on the other hand accounts for the uncertainties in prediction by estimating output distributions instead of rigid labels [12]. The Dempster-Shafer theory of evidence (DST) [26] and Subjective Logic (SL) provide a well-defined approach to quantifying class uncertainties. The approach suggests that for a $K$ class network, with a belief mass $b_{k}$ assigned to each class label $k\in K$, an overall uncertainty is $u$ can be expressed, such that:
@@ 2.2 Evidential Deep Learning
The output of the neural network model (before the conventional SoftMax layer, referred to as logits) is trained to represent the evidence vectors $e_{k}$. These evidence vectors represent the amount of support on how confident a sample is, given a particular class. The vectors relate to the belief masses and uncertainties as follows:
@@
It is interesting to note that the uncertainty value from Equation 4 is inversely proportional to $S$, which represents the sum of evidence and the number of labels. This notes that a zero-evidence vector would yield a belief mass $b_{k}=0$, and the maximum uncertainty $u=1$. The likelihood function for classification models learns over a categorical distribution. EDL places a higher order distribution over the parameters of the categorical distribution (label probabilities) to allow for prediction uncertainty estimation $u$. To ensure a tractable posterior, it is essential to select a prior conjugate distribution, such as the Dirichlet distribution. Parameterized by $\alpha=[\alpha_{1},...,\alpha_{K}]$, the Dirichlet distribution’s density function is expressed by:
@@
The $S$ here is taken from Equation LABEL:uncertainty and can be referred to as the strength of the Dirichlet distribution.
@@ 2.3 Conformal Prediction
Figure 1: Conformal Prediction Framework. The dataset is split into a training set, a calibration set, and a testing set. Conformal Prediction utilizes the scoring function and calibration set to generate a quantile value. At inference, a test sample is passed to the train network, and the output score along with the quantile value is used to generate prediction sets.
3 Expirements and Results
3.1 Dataset
This paper focuses on the challenge of classifying different types of skin lesions by leveraging deep learning approaches and medical images. For this purpose, we work on the HAM10000 dataset [27], which contains more than 10,000 images captured from real patients with seven distinct lesion classes. The HAM10000 dataset is highly imbalanced, with some classes having more samples than others. To mitigate bias, we applied data augmentation techniques such as flipping and rotations to increase the number of samples in less-represented classes. To train the network, we divided the dataset into 50% training, 20% validation, and 30% for testing. We used samples from both validation and testing datasets to form the calibration dataset used in CP.
3.2 Expiremental Setup
3.2.1 Model Specifications and Training configuration
In all our experiments, ResNet-18 [28] was adopted as the underlying core network for all the uncertainty quantification methods. The network consists of 18 layers, which include convolutional layers, batch normalization, dropout layers, ReLU activation functions, and residual blocks. A classification head is modified to classify the input images into their corresponding 7 classes. We trained the core model using the cross-entropy loss for 100 epochs. The Adam optimizer was used with a learning rate of 1e-4. A learning rate scheduler was added to reduce the learning rate by 0.5 when the validation loss is not improved for 10 consecutive epochs. All the experiments are run on an NVIDIA RTX A5000.
3.2.2 Uncertainty Quantification Models
As previously indicated, the ResNet-18 model serves as the foundational architecture for implementing uncertainty quantification methods. However, distinct modifications are necessitated to incorporate each of the three uncertainty quantification techniques. In this section, we elaborate on the required design settings tailored for the implementation of each method. Furthermore, we address any supplementary training configurations and methodologies employed in the integration of the core model to the uncertainty quantification technique.
@@ 3.3 Evaluation of Uncertainty Quantification methods
This section presents a quantitative assessment and comparative analysis of the three uncertainty quantification methods: CP, MCD, and EDL. Initially, the base classification model was trained without the incorporation of any uncertainty method. The model achieves an 86.7% accuracy rate on the testing dataset that contains 1103 samples. In the context of CP, we set a user-defined confidence interval of $1-\alpha$, where $\alpha$ is chosen to be 0.10, corresponding to a 90% confidence level. This implies that the ground truth label is encompassed within the generated prediction sets with a 90% confidence bound. A larger prediction set indicates higher predictive uncertainty, while a smaller set implies lower uncertainty. As mentioned earlier, CP operates as an add-on algorithm, acting on the SoftMax probabilities produced by the core model. Consequently, deploying CP does not alter the initial accuracy of the model. In the context of predictive uncertainty, CP assigns a mean uncertainty value of 0.40 to correctly classified samples and 0.79 to wrongly classified samples. Meanwhile, for the MCD method, 1000 forward passes are executed during inference to estimate the model’s uncertainty, resulting in an accuracy of 85.8%. Noteworthy is that correctly classified examples in MCD are associated with a mean uncertainty value of 0.01, whereas wrongly classified examples exhibit a mean uncertainty value of 0.09. In contrast, EDL takes a distinct approach by being trained from scratch using the mean square error loss that explicitly accounts for uncertainties. The EDL model achieves an accuracy of 85.7%, with a mean uncertainty value of 0.19 assigned to correctly classified samples and 0.51 to wrongly classified examples. The results are shown in Table 1.
UQ algorithm Accuracy $\mathbf{U_{\text{correct}}\pm\text{std}}$ $\mathbf{U_{\text{wrong}}\pm\text{std}}$ Retraining
@@
CP 86.7% $0.40\pm 0.3$ $0.79\pm 0.15$
Table 1: Comparison of UQ Algorithms: Accuracy, Uncertainty Estimates, and Retraining Indicator. The table presents model accuracy along with mean uncertainty values for both correctly classified samples and wrongly classified samples, in addition to an indication of whether retraining was performed, for Monte Carlo Dropout (MCD), Evidential Deep Learning (EDL), and Conformal Prediction (CP).
Figure 2 showcases the histogram distributions of the assigned uncertainty values for correctly and wrongly classified samples, across the three aforementioned uncertainty quantification methods. The same is also illustrated in Figure 3 through the kernel density estimation plot, which offers a continuous representation of the probability density function for a more comprehensive understanding of the uncertainty characteristics across the dataset.
@@
Figure 2: The histogram distribution of uncertainty values assigned by each uncertainty quantification algorithm to both correctly and wrongly classified samples.
(a) Correctly classified samples
(b) Wrongly classified samples
Figure 3: The density distribution of uncertainty values assigned by each uncertainty quantification algorithm to both correctly and wrongly classified samples.
3.3.1 Discussion
In light of the conducted experiments, the obtained results provide valuable insights and analysis into uncertainty prediction techniques. Before discussing the advantages and drawbacks of each technique, it is imperative to emphasize that predictive uncertainty should not be misconstrued as a definitive indicator of classification output. Instead, it serves as supplementary information that enriches the model’s robustness. It’s typical to find correctly classified samples in the dataset with a mix of both high and low uncertainty values. The variability in uncertainty values is a result of the differences in difficulty levels among individual samples. Ideally, the uncertainty value of the wrongly classified samples should be consistently higher, indicating less confidence in the network’s output. As observed in Figure 2 and Figure 3, CP exhibits such behavior, where the majority of wrongly classified samples are assigned high uncertainty values. When examining the results for both MCD and EDL, we observe a flatter distribution for wrongly classified samples, suggesting a greater variability in uncertainty values.
Compared with MCD and EDL, CP stands out for its simplicity in implementation. The algorithm operates seamlessly, eliminating the need for retraining procedures or modifications to the model’s architecture. This crucial feature facilitates the quantification of uncertainties and ensures that the accuracy of the original model remains unaffected. In MCD, little modifications to the model architecture might be necessary to include dropout layers if not present. Nevertheless, MCD requires running multiple forward passes, contributing to an increased computational load. Moreover, it is crucial to note that high dropout rates have the potential to compromise accuracy, emphasizing the importance of carefully tuning dropout parameters to strike a balance between uncertainty quantification and model performance. Additionally, some studies show that MCD tends to output ill-calibrated uncertainty values that can be unreliable for uncertainty quantification [22]. EDL, on the other hand, generates sharp uncertainty distribution for the correctly classified samples, but a broad distribution with wrongly classified samples, making it less reliable. The implementation of EDL is relatively more complex when compared to CP and MCD. It requires modifying the core model, changing the training loss function, and retraining the modified model from scratch.

### 16 Mar 2024 v2

@@ Abstract
Background and objective: Uncertainty quantification is a pivotal field that contributes to realizing reliable and robust systems. It becomes instrumental in fortifying safe decisions by providing complementary information, particularly within high-risk applications. existing studies have explored various methods that often operate under specific assumptions or necessitate substantial modifications to the network architecture to effectively account for uncertainties. The objective of this paper is to study Conformal Prediction, an emerging distribution-free uncertainty quantification technique, and provide a comprehensive understanding of the advantages and limitations inherent in various methods within the medical imaging field.
Methods: In this study, we developed Conformal Prediction, Monte Carlo Dropout, and Evidential Deep Learning approaches to assess uncertainty quantification in deep neural networks. The effectiveness of these methods is evaluated using three public medical imaging datasets focused on detecting pigmented skin lesions and blood cell types.
Results: The experimental results demonstrate a significant enhancement in uncertainty quantification with the utilization of the Conformal Prediction method, surpassing the performance of the other two methods. Furthermore, the results present insights into the effectiveness of each uncertainty method in handling Out-of-Distribution samples from domain-shifted datasets. Our code is available at:
Conclusions: Our conclusion highlights a robust and consistent performance of conformal prediction across diverse testing conditions. This positions it as the preferred choice for decision-making in safety-critical applications.
@@ 1 Introduction
In medical applications, deep learning plays a vital role by enabling more accurate and efficient analysis of complex data, including medical image classification, tumor detection, and medical image segmentation. Deep learning models, however, are trained on clean and well-curated datasets, which enables models to achieve high accuracies on such data. However, images are often subjected to noise, artifacts, and inherent limitations of the imaging process. These factors can potentially impact the predictions of deep learning models, leading to uncertainty in their outputs. Therefore, it is crucial to express confidence in the network’s predictions to assess the reliability and robustness of its results, especially in critical applications such as medical imaging.
@@
In machine learning, two sources of uncertainties collectively influence model predictions [7]. The first source, known as epistemic uncertainty, arises from a lack of knowledge and understanding. Model retraining and incorporating additional information and data are typically employed to mitigate this type of uncertainty [8]. The second source of uncertainty is aleatoric uncertainty, which characterizes the inherent randomness within the data and, as such, remains irreducible. In the existing literature, various approaches offer both theoretical and practical solutions for quantifying uncertainties associated with deep learning models [9, 10]. These methods can be broadly categorized into Single Deterministic Networks, Bayesian Neural Networks, Deep Ensembles, and Test-time Augmentation methods [11].
@@
Single deterministic networks are straightforward methods to represent the predictive uncertainty of deep learning models. As suggested by the name, the networks are trained to adjust their weights in a deterministic manner, and the uncertainty is measured through a single forward pass. A simple baseline of single deterministic networks is the utilization of the neural network output probabilities (i.e., SoftMax outputs for classification) as an interpretation of the model uncertainty. Several studies, however, showed that neural networks are often over-confident [12], hence their output probabilities are miscalibrated [13] and tend to generate wrong uncertainty estimates. A more generic approach in single deterministic networks includes predicting the parameters of a prior distribution over the outputs. such algorithms are known as Evidential learning [14] or Prior Networks [15]. The framework of these approaches involves introducing a higher-order distribution over the likelihood function, with the parameters of these distributions being estimated using the underlying neural network. The predicted class probability corresponds to the expected value of the distribution, while the variance of the distribution serves as an estimate of the model’s uncertainty. In [16], Feng et al. proposed a vision transformer-based backbone as a reliable multi-scale classification framework for Whole Slide Images. The backbone is integrated with evidence uncertainty theory to assess uncertainty levels at different magnifications of a microscope. Despite the demonstrated outperforming results, evidence-based approaches require changes to the model architecture and the loss function employed at training.
@@
Bayesian Neural Networks (BNNs) [17] seamlessly merge the principles of Bayesian inference with traditional neural networks. Instead of considering BNN parameters as fixed, deterministic values, they are treated as probability distributions. This unique approach empowers BNNs to effectively capture and represent predictive uncertainty. Consider a neural network with parameters $\omega$ and a training dataset $D_{t}=(x,y)$. If the network weights are assigned a prior distribution, denoted as $P(\omega)$, the posterior distribution $P(\omega|x,y)$ can be calculated using Bayes’ theorem. To calculate the posterior $p(\omega|X,Y)$, it necessitates integrating over all possible values of $\omega$, a task that is often intractable. Generally, obtaining the posterior is only viable through either Variational Inference [18], where a simple and tractable distribution is used to approximate the exact posterior, or through other approximation techniques, such as Markov Chain Monte Carlo (MCMC) [19]. Leibig et al. [20] utilized MCDropout to assess the uncertainty of deep learning models in detecting diabetic retinopathy. while the method captures useful uncertainty information, the sampling process is computationally expensive, hence it hinders real-time implementation.
@@
Recently, Conformal prediction (CP) [22, 23] emerged as a prominent uncertainty quantification technique. It has gained attention in the fields of computer vision [24, 25], and more specifically, in medical imaging [26]. In addition to its elegance and computational efficiency, CP algorithms belong to the domain of ’distribution-free’ methodologies. This implies that they are model-agnostic, data distribution agnostic, and applicable to finite sample sizes.
In this paper, we compare three top-performing uncertainty quantification methods, namely CP, Monte Carlo Dropout (MCD), and Evidential Deep Learning (EDL), and provide a comprehensive and detailed analysis of their performance. The selection of these three methods is attributed to the fact that they cover diverse approaches, including Bayesian, evidential, and distribution-free methodologies. Furthermore, our selection process of the algorithms encompassed a range of implementation details, spanning from maintaining the original network structure without modification to undertaking complete retraining and adjusting the loss functions. The central research questions under examination are: How stable and robust are the prominent uncertainty quantification methods, and how do they perform individually in real environment settings? We include three experiments to underscore the unique attributes of each method, with a particular emphasis on CP, a less-explored approach for skin lesion classification tasks. The main contributions of our work are as follows:
@@
2 Methods
2.1 Dataset
This paper focuses on the challenge of classifying different types of skin lesions by leveraging deep-learning approaches and medical images. For this purpose, we work on three publicly available medical datasets, two skin lesion datasets named: the HAM10000 dataset [27], and the Dermofit (DMF) and Blood Cell Microscope (BCM) [acevedo2020dataset] dataset. [28]. The HAM10000 dataset contains more than 10,000 images and the DMF contains more than 1200 images, both captured from real patients with seven distinct lesion classes. Both datasets are highly imbalanced, with some classes having more samples than others. To address biases stemming from imbalanced classes, we initially identified the class with the highest sample count. Utilizing this count as a benchmark, we calculated the additional samples required for each class to achieve balance. Subsequently, various data augmentation methods, including vertical and horizontal flipping, angle rotations, and cropping, were employed to generate the necessary number of samples for each class. The BCM dataset contains around 17,000 peripheral blood cell images labeled with 8 different blood cell samples. To train the network, we divided the dataset into 50% training, 20% validation, and 30% for testing. For CP, we reserved a portion of samples from both validation and testing datasets that were not used before and utilized them for calibration.
2.2 Model Specifications and Training configuration
In all our main experiments, we employ ResNet-18 [29] as the underlying core network for all uncertainty quantification methods. The network, which is pretrained on ImageNet, consists of 18 layers that include convolutional layers, batch normalization, dropout layers, ReLU activation functions, and residual blocks. In addition, we have also utilized two additional networks: ResNet-50 and VGG-11 to study the effect of changing the core network. In all cases, a classification head is modified to classify the input images into their corresponding classes. We trained the core model using the cross-entropy loss for 100 epochs. The Adam optimizer was used with a learning rate of 1e-4. A learning rate scheduler was added to reduce the learning rate by 0.5 when the validation loss is not improved for 10 consecutive epochs. All the experiments are run on an NVIDIA RTX A5000.
2.3 Uncertainty Quantification Methods
In the rich domain of uncertainty quantification, understanding the nuances of existing methodologies and their implementation details is paramount. This section delves into the background literature and related work that forms the foundation for uncertainty methods. We specifically focus on three quantification techniques that form the basis of our experimental analysis. For a more comprehensive review of uncertainty quantification methods in deep learning, interested readers can refer to [11, 30].
2.3.1 Monte Carlo Dropout
@@
Let $f(x)$ be the prediction of the neural network for input $x$. During MCD, we obtain $T$ samples of predictions $f_{t}(x)$ where $t=1,2,...,T$ represents each Monte Carlo sample. These predictions are obtained by performing $T$ forward passes through the network. The mean prediction can be represented as:
@@
2.3.2 Evidential Deep Learning
In a traditional Convolution Neural Network (CNN), label probabilities are generated by applying a SoftMax activation function to the last layer. The objective is to maximize the likelihood function by minimizing a metric of loss, often a cross-entropy loss, which tends to make overconfident predictions [32]. The output of the SoftMax operation reflects discrete point estimates of class probabilities without proper indication of the uncertainty present in the estimate. EDL on the other hand accounts for the uncertainties in prediction by estimating output distributions instead of rigid labels [14]. The intuition behind EDL lies in modeling the network predictions as a distribution over the categorical SoftMax outputs. Sampling from this distribution yields the realization of the uncertainty through the variance of the distribution.
The Dempster-Shafer theory of evidence (DST) [33] and Subjective Logic (SL) provides a well-defined framework for quantifying class uncertainties. The approach suggests assigning a belief mass $b_{k}$ for each of the $K$ classes of the network. $b_{k}$ represents the belief that the truth can be any of the possible states. Hence the overall uncertainty $u$ can be expressed by computing the residual sum of all belief masses:
@@ 2.3.2 Evidential Deep Learning
In EDL, The SoftMax layer is removed, and the neural network is trained using a specific loss function to output evidence vectors $e_{k}$. These evidence vectors represent the amount of support on how confident a sample is, given a particular class. The vectors relate to the belief masses and uncertainties as follows:
@@
It is interesting to note that the uncertainty value from Equation 4 is inversely proportional to $S$, which represents the sum of evidence and the number of labels. This notes that a zero-evidence vector would yield a belief mass $b_{k}=0$, and the maximum uncertainty $u=1$.
The likelihood function for classification models learns over a categorical distribution. EDL places a higher order distribution over the parameters of the categorical distribution (label probabilities) to allow for prediction uncertainty estimation $u$. To ensure a tractable posterior, it is essential to select a prior conjugate distribution, such as the Dirichlet distribution. Parameterized by $\alpha=[\alpha_{1},...,\alpha_{K}]$, the Dirichlet distribution’s density function is expressed by:
@@
The $S$ here is taken from Equation 4 and can be referred to as the strength of the Dirichlet distribution.
@@ 2.3.3 Conformal Prediction
Figure 1: Conformal Prediction Framework. The dataset is split into a training set, a calibration set, and a testing set. Conformal Prediction utilizes the scoring function and calibration set to generate a quantile value $\hat{q}$. At inference, a test sample is passed to the train network, and the output score along with the quantile value is used to generate prediction sets.
2.4 Uncertainty Quantification Settings
As previously indicated, ResNet-18 serves as the foundational architecture for implementing uncertainty quantification methods. However, distinct modifications are necessitated to incorporate each of the three uncertainty quantification techniques. In this section, we elaborate on the required design settings tailored for the implementation of each method. Furthermore, we address any supplementary training configurations and methodologies employed in the integration of the core model to the uncertainty quantification technique. It’s crucial to note that due to the differing implementation procedures required by each uncertainty quantification technique—including distinct training approaches and loss functions—each model will ultimately exhibit varying classification accuracies, subject to the training of the networks.
@@ 3.1 Evaluation of Uncertainty Quantification methods
This section presents a quantitative assessment and comparative analysis of the three uncertainty quantification methods: CP, MCD, and EDL. In this study, our investigation relies on examining the distribution of quantified uncertainty values assigned by each approach. To assess performance comprehensively, we utilized both the mean uncertainty value of correctly classified samples and that of wrongly classified samples. A consistent and robust uncertainty quantification approach lies in its ability to assign higher uncertainty values to wrongly classified samples compared to correctly classified samples.
Initially, the base classification model was trained without the incorporation of any uncertainty method. The model achieves an 86.7% accuracy rate on the HAM10000 testing dataset that contains 1103 samples. In the context of CP, we set a user-defined confidence interval of $1-\alpha$, where $\alpha$ is chosen to be 0.10, corresponding to a 90% confidence level. This implies that the ground truth label is encompassed within the generated prediction sets with a 90% confidence bound. A larger prediction set indicates higher predictive uncertainty, while a smaller set implies lower uncertainty. As mentioned earlier, CP operates as an add-on algorithm, acting on the SoftMax probabilities produced by the core model. Consequently, deploying CP does not alter the initial accuracy of the model. In the context of predictive uncertainty, the average uncertainty value of all correctly classified samples is 0.4, while the average uncertainty value of all wrongly classified samples is 0.79.
Meanwhile, for the MCD method, 1000 forward passes are executed during inference to estimate the model’s uncertainty, resulting in accuracies of 85.8%. Noteworthy is that correctly classified examples in MCD are associated with a mean uncertainty value of 0.01, whereas wrongly classified examples exhibit a mean uncertainty value of 0.09. In contrast, EDL takes a distinct approach by being trained from scratch using the mean square error loss that explicitly accounts for uncertainties. The EDL model achieves an accuracy of 85.7%, with a mean uncertainty value of 0.19 assigned to correctly classified samples and 0.51 to wrongly classified examples.
Additionally, we have conducted a similar analysis of the three uncertainty quantification methods on the DMF dataset and the BCM dataset. The detailed results are shown in Table 1. Moreover, the histogram distributions of the assigned uncertainty values for correctly and wrongly classified samples, across the three aforementioned uncertainty quantification methods are shown in Figure 2.
UQ method Accuracy $\mathbf{U_{\text{correct}}\pm\text{std}}$ $\mathbf{U_{\text{wrong}}\pm\text{std}}$
HAM10000
@@
CP 86.7% $0.40\pm 0.30$ $0.79\pm 0.15$
DMF
MCD 75.2% $0.05\pm 0.03$ $0.08\pm 0.02$
EDL 70.4% $0.24\pm 0.17$ $0.47\pm 0.11$
CP 72.7% $0.33\pm 0.25$ $0.57\pm 0.18$
BCM
MCD 98.0% $0.01\pm 0.02$ $0.07\pm 0.03$
EDL 98.2% $0.42\pm 0.06$ $0.48\pm 0.06$
CP 98.6% $0.34\pm 0.29$ $0.78\pm 0.16$
Table 1: Comparison of UQ Algorithms: Accuracy, Average Uncertainty Estimates for HAM10000, DMF, and BCM datasets. The table presents model accuracy along with average uncertainty values for both correctly classified samples and wrongly classified samples, as determined by the core network and per their respective ground truth labels. Algorithms include Monte Carlo Dropout (MCD), Evidential Deep Learning (EDL), and Conformal Prediction (CP).
@@
Figure 2: The histogram distribution of uncertainty values assigned by each uncertainty quantification algorithm to both correctly and wrongly classified samples. Results illustrate that while both MCD and EDL assign diverse values of uncertainties to wrongly classified samples, CP, on the other hand, assigns high uncertainty values to those samples.
Next in our study, we extended our analysis and experimental investigations by including two additional models: ResNet-50 and VGG-11. The objective is to evaluate how altering the core network impacts the uncertainty results generated by each of the employed methods. The structure of ResNet-50 shares similarities with ResNet-18, albeit being deeper with over 50 convolutional layers. VGG-11 on the other hand, consists of multiple layers of convolutional and max-pooling operations, followed by fully connected layers. Despite its straightforward design, VGG-11 has demonstrated competitive performance across a wide range of image classification tasks. The same training procedures that were implemented in Section 2.2 are followed in these experiments. The ResNet-50 model achieved similar results to the initial experiments. We observe that there is no significant impact on uncertainty values generated by using ResNet-50. In contrast, VGG-11 network did not achieve high classification accuracies, and hence the worse performance of the CP uncertainty quantification. We record the accuracies and average uncertainties assigned by each method in Table 2. Moreover, the distribution of the uncertainty values assigned for both the correctly classified examples and the wrongly classified examples for the three uncertainty methods are shown in Figure 3
UQ algorithm ResNet-50 VGG-11
Accuracy $\mathbf{U_{\text{correct}}\pm\text{std}}$ $\mathbf{U_{\text{wrong}}\pm\text{std}}$ Accuracy $\mathbf{U_{\text{correct}}\pm\text{std}}$ $\mathbf{U_{\text{wrong}}\pm\text{std}}$
MCD 88.1% $0.01\pm 0.03$ $0.09\pm 0.05$ 83.7% $0.01\pm 0.02$ $0.05\pm 0.04$
EDL 87.4% $0.21\pm 0.22$ $0.73\pm 0.29$ 80.2% $0.30\pm 0.36$ $0.87\pm 0.27$
CP 88.0% $0.46\pm 0.28$ $0.75\pm 0.18$ 81.0% $0.35\pm 0.28$ $0.61\pm 0.23$
Table 2: Comparison of UQ Algorithms: Accuracy, average Uncertainty Estimates for ResNet-50 and VGG-11. The table presents model accuracy along with average uncertainty values for both correctly classified samples and wrongly classified samples, as determined by the core network and per their respective ground truth labels. Algorithms include Monte Carlo Dropout (MCD), Evidential Deep Learning (EDL), and Conformal Prediction (CP).
(a) Correctly classified samples: ResNet-50
(b) Wrongly classified samples: ResNet-50
(c) Correctly classified samples: VGG-11
(d) Wrongly classified samples: VGG-11
Figure 3: Histogram of uncertainty values assigned to correctly classified samples and wrongly classified samples using 2 different deep learning models: ResNet-50 and VGG-11. The three uncertainty quantification methods are MCDropout, EDL, and CP
Finally, we conducted further analysis by examining samples from the HAM10000 dataset along with their corresponding uncertainty values, categorized into distinct groups based on their classification outcomes by the core network. These groups were further stratified into correctly classified and wrongly classified samples, with their uncertainty levels quantified using the conformal prediction algorithm. The samples are shown in Figure 4. Notably, a discernible difference in skin tone is observed between misclassified and correctly classified samples. Such potential bias in the dataset can adversely affect network accuracy. Furthermore, we observed that misclassified samples with low uncertainty exhibited more prominent skin lesion patches compared to those with high uncertainty. This disparity in patch visibility could account for the lower uncertainty associated with the former group.
Figure 4: Analysis of samples with corresponding uncertainty values. Samples are categorized into correctly classified and wrongly classified groups, with their uncertainty quantified using the conformal prediction algorithm. Misclassified samples often exhibit a noticeably different skin tone compared to correctly classified ones. Additionally, misclassified samples with low uncertainty display more visible skin lesion patches compared to those with high uncertainty, possibly influencing the model’s predictive confidence.